# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ Appendix C Theoretical Material
where, for each pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ the sum increases by one if the condition of Eq. 43 is satisfied.
Among these optimal pairs, there are those possessing the intended semantics by Lemma C.4, which consists of possible permutations of the concepts and element-wise invertible transformations of the concept values. We start by evaluating the number of possible element-wise invertible transformations for each subset of concepts $\mathcal{G}_{i}\subset\mathcal{G}$.
Each concept $c_{i}$ has values in the subset $\mathcal{G}_{i}$, for a total of $|\mathcal{G}_{i}|$ values. The overall number of possible invertible transformations is then given by the number of possible permutations, resulting in a total of $|\mathcal{G}_{i}|!$ maps. This means that the total number of element-wise invertible transformations is given by:
@@
where $m(\bullet)$ counts how many repetitions are present. When the multi-set contains repeated numbers, i.e., when different $c_{i}$ and $c_{j}$ have the same cardinality, we have to account a possible permutation of the concept indices. Therefore, for each element $\xi\in s\mathcal{G}$ we have that the total number of permutations of concepts amount to:
@@
Finally, the total number of pairs possessing the intended semantics amount to $\mathrm{perm}(\mathcal{G})\times\prod_{i=1}^{k}|\mathcal{G}_{i}|!$, resulting that the total number of deterministic JRSs are:
@@
The proof follows immediately by replacing $\mathsf{Vert}(\mathcal{B})$ with $\{\bm{\mathrm{\beta}}^{*}\}$, allowing only for the ground-truth inference function. In this context, by following similar steps of Theorem 3.5, we arrive at a similar count to Eq. 44, that is:
@@
In the context of RSs, the only pair possessing the intended semantics is $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$, which has to subtracted from Eq. 49. This shows the claim. ∎
@@
Consider a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$, where both $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$ can be non-deterministic conditional probability distributions. Step (i). We begin by first recalling Lemma C.1. Under Assumptions 3.1 and 3.2, a CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ is optimal, i.e., it attains maximum likelihood, if it holds that:
@@
Notice that, according to Assumptions 3.1 and 3.2, it also holds that the labels must be predicted with probability one to attain maximum likelihood, that is:
@@
Now consider a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$. By Lemma C.1 (i), it holds that to obtain maximum likelihood, the learned knowledge $\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$ must be deterministic:
@@
which implies that, necessarily, to have optimal $\bm{\mathrm{\beta}}$ the space $\mathcal{B}$ restricts to the set $\mathsf{Vert}(\mathcal{B})$, so any learned knowledge $\bm{\mathrm{\beta}}\in\mathsf{Vert}(\mathcal{B})$. This shows that optimal pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ are in $\mathcal{A}\times\mathsf{Vert}(\mathcal{B})$.
Step (ii). In this step, we make use of Assumption 3.7 and consider only inference functions $\bm{\mathrm{\omega}}$ with that property. We start by considering a couple $(\bm{\mathrm{a}}^{\prime},\bm{\mathrm{b}}^{\prime})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ that attain maximum likelihood, that is according to Theorem 3.5:
@@
Now, we have to prove that, given $\bm{\mathrm{b}}^{\prime}\in\mathsf{Vert}(\mathcal{B})$, there are no $\bm{\mathrm{\alpha}}\in\mathcal{A}\setminus\mathsf{Vert}(\mathcal{A})$ that can attain maximum likelihood for at least one possible choice of $\bm{\mathrm{\omega}}^{\prime}\in\Omega$, such that for all $\bm{\mathrm{c}}\in\mathcal{C}$ it holds $\bm{\mathrm{\omega}}^{\prime}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})=\bm{\mathrm{b}}^{\prime}(\bm{\mathrm{c}})$. Since $\mathcal{A}$ is a simplex, we can always construct a non-deterministic conditional probability distribution $\bm{\mathrm{\alpha}}\in\mathcal{A}$ from a convex combination of the vertices $\bm{\mathrm{a}}_{i}\in\mathsf{Vert}(\mathcal{A})$, i.e., for all $\bm{\mathrm{g}}\in\mathcal{G}$ it holds:
@@
where $\lambda\in(0,1)$. When the deterministic JRSs count (Eq. 7) equals to zero, we know that by Theorem 3.5 only $(\bm{\mathrm{a}}^{\prime},\bm{\mathrm{b}}^{\prime})$ attains maximum likelihood, whereas $(\bm{\mathrm{a}}^{\prime\prime},\bm{\mathrm{b}}^{\prime})$ is not optimal. Therefore, there exists at least one ^$\hat{}\bm{\mathrm{g}}\in\mathsf{supp}(p^{*}(\bm{\mathrm{G}}))$ such that
@@
We now have to look at the form of $\bm{\mathrm{f}}$ that can induce such a non-deterministic $\bm{\mathrm{\alpha}}=\lambda\bm{\mathrm{a}}^{\prime}+(1-\lambda)\bm{\mathrm{a}}^{\prime\prime}$. Recalling, the definition of $\bm{\mathrm{\alpha}}$ (Eq. 2) we have that:
@@
(2) The remaining option is that $\bm{\mathrm{f}}:\mathbb{R}^{n}\to\Delta_{\mathcal{C}}$, excluding mapping to the vertices almost everywhere in $\mathsf{supp}(p^{*}(\bm{\mathrm{X}}\mid\hat{}\bm{\mathrm{g}}))$. In light of this, we rewrite $\bm{\mathrm{f}}$ as follows:
@@
where $p(\bm{\mathrm{c}}\mid\bm{\mathrm{x}}):=\bm{\mathrm{f}}(\bm{\mathrm{x}})_{\bm{\mathrm{c}}}$. Let ^$\hat{}\bm{\mathrm{c}}^{\prime}:=\bm{\mathrm{a}}^{\prime}(\hat{}\bm{\mathrm{g}})$ and ^$\hat{}\bm{\mathrm{c}}^{\prime\prime}:=\bm{\mathrm{a}}^{\prime\prime}(\hat{}\bm{\mathrm{g}})$. But $\bm{\mathrm{\alpha}}$ is a convex combination of $\bm{\mathrm{a}}^{\prime}$ and $\bm{\mathrm{a}}^{\prime\prime}$, hence the function $\bm{\mathrm{f}}$ must attribute non-zero probability mass only to the concept vectors ^$\hat{}\bm{\mathrm{c}}^{\prime}$ and ^$\hat{}\bm{\mathrm{c}}^{\prime\prime}$. Hence, we rewrite it as:
@@
where $\mathcal{X}$ has measure one. Notice that in this case it holds $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{b}}^{\prime}(\hat{}\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\neq\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{b}}^{\prime}(\hat{}\bm{\mathrm{c}}^{\prime\prime})_{\bm{\mathrm{y}}}$. By Assumption 3.7 we have that, for any choice of $\lambda\in(0,1)$ it holds that for all $\bm{\mathrm{x}}\in\mathcal{X}$:
@@
where the last step in Eq. 61 follows from the fact that $\bm{\mathrm{b}}^{\prime}\in\mathsf{Vert}(\mathcal{B})$, so giving point-mass conditional probability distributions.
(1) and (2) together show that any $\bm{\mathrm{\alpha}}\in\mathcal{A}\setminus\mathsf{Vert}(\mathcal{A})$ constructed as a convex combination of $\bm{\mathrm{a}}^{\prime}$ and an $\bm{\mathrm{a}}^{\prime\prime}\neq\bm{\mathrm{a}}^{\prime}$ cannot attain maximum likelihood. Hence, the optimal pairs restrict to $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ and since the count of JRSs (Eq. 7) is zero, all $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ that are optimal also possess the intended semantics. This concludes the proof. ∎
@@ Appendix D Impact of mitigation strategies on the deterministic JRSs count
We focus on how the count can be updated for some mitigation strategies that we accounted for in Section 4. We consider those strategies that gives a constraint for the Eq. 7, namely: multi-task learning, concept supervision, knowledge distillation, and reconstruction.
@@ D.1 Multi-task learning
Suppose that for a subset $\mathcal{G}^{\tau}\subseteq\mathrm{supp}(p^{*}(\bm{\mathrm{G}}))$, input examples are complemented with additional labels $\bm{\mathrm{y}}^{\tau}$, obtained by applying an inference layer $\bm{\mathrm{\beta}}_{\mathsf{K}^{\tau}}(\bm{\mathrm{g}})$. Here, $\mathsf{K}^{\tau}$ represents the augmented knowledge, giving a set $\mathcal{Y}^{\tau}$ of additional labels. To understand how such a choice modifies the count of jrss, we have to consider those $\bm{\mathrm{\beta}}:\mathcal{C}\to\mathcal{Y}\times\mathcal{Y}^{\tau}$, where a component $\bm{\mathrm{\beta}}_{\mathcal{Y}}$ maps to the original task labels $\bm{\mathrm{Y}}$ and the other component $\bm{\mathrm{\beta}}_{\mathcal{Y}^{\tau}}$ maps to the augmented labels $\bm{\mathrm{Y}}^{\tau}$. Let :=×$\mathcal{B}^{*}:={\mathcal{B}}\times{\mathcal{B}^{\tau}}$ the space where such functions are defined. As obtained in [Marconato et al., 2023b], the constraint for a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B}^{*})$ can be written as:
@@
This term can be steadily combined with the one appearing in Theorem 3.5 to yield a reduction of possible maps $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ that successfully allow a function $\bm{\mathrm{\beta}}\in\mathcal{B}^{*}$ to predict both $(\bm{\mathrm{y}},\bm{\mathrm{y}}^{\tau})$ consistently. Notice that, in the limit where the augmented knowledge comprehends enough additional labels and the support covers the whole , the only admitted solutions consist of:
@@ D.2 Concept supervision
The count reduction was studied in [Marconato et al., 2023b], and here we report the same term. Consider the case where we supply concept supervision for a subset of concepts $\bm{\mathrm{G}}_{\mathcal{I}}\subseteq\bm{\mathrm{G}}$, with $\mathcal{I}\subseteq[k]$, and then augmenting the log-likelihood objective with a cross-entropy loss over the concepts: $\sum_{i\in\mathcal{I}}\log p_{\theta}(C_{i}=g_{i}\mid\bm{\mathrm{x}})$. We consider also a subset $\mathcal{G}^{C}\subseteq\mathsf{supp}(p^{*}(\bm{\mathrm{G}}))$ where this supervision is provided. The constraint on the number of deterministic $\bm{\mathrm{\alpha}}$’s is given by the following
@@
If $\mathcal{I}=[k]$ and $\mathcal{G}^{C}=\mathcal{G}$, then we get $\bm{\mathrm{\alpha}}\equiv\mathrm{id}$. Naturally, this requires dense annotations for all possible inputs, which may be prohibitive.

### v2

@@ Appendix C Theoretical Material
where, for each pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ the sum increases by one if the condition of Eq.˜45 is satisfied.
Among these optimal pairs, some possess the intended semantics and therefore do not count as JRSs. By Lemma˜C.4, these correspond to all possible permutations of the concepts combined with all possible element-wise invertible transformations of the concept values.
We begin by evaluating the number of possible element-wise invertible transformations. The $i$-th concept can attain $|\mathcal{G}_{i}|$ values. The overall number of possible invertible transformations is then given by the number of possible permutations, resulting in a total of $|\mathcal{G}_{i}|!$ maps. Hence, the number of element-wise invertible transformations is:
@@
where $m(\bullet)$ counts how many repetitions are present. When different $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ have the same cardinality, we can permute $G_{i}$ and $G_{j}$ without compromising optimality, as there always exists a $\bm{\mathrm{\beta}}$ that inverts the permutation and thus provides the same output. Thankfully, this shows in the multiplicity $m(\cdot)$ and we can account for this. Therefore, for each element $\xi\in s\mathcal{G}$ we have that the total number of permutations of concepts amount to:
@@
Putting everything together, we obtain that the number of optimal pairs possessing the intended semantics is $\mathrm{perm}(\mathcal{G})\times\prod_{i=1}^{k}|\mathcal{G}_{i}|!$, meaning that the total number of deterministic JRSs is:
@@
The proof follows immediately by replacing $\mathsf{Vert}(\mathcal{B})$ with $\{\bm{\mathrm{\beta}}^{*}\}$, allowing only for the ground-truth inference function. In this context, by following similar steps of Theorem˜3.6, we arrive at a similar count to Eq.˜46, that is:
@@
In the context of regular RSs, the only pair possessing the intended semantics is $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$, which has to subtracted from Eq.˜51. This proves the claim. ∎
@@
Consider a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$, where both $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$ can be non-deterministic conditional probability distributions.
Step (i). We begin by first recalling Lemma˜C.1. Under Assumptions˜3.1 and 3.2, a CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ is optimal, i.e., it attains maximum likelihood, if it holds that:
@@
Notice that, according to Assumptions˜3.1 and 3.2, it also holds that the labels must be predicted with probability one to attain maximum likelihood, that is:
@@
Now consider a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$. By Lemma˜C.1 (i), it holds that to obtain maximum likelihood, the learned knowledge $\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$ must be deterministic:
@@
which implies that, necessarily, for $\bm{\mathrm{\beta}}$ to be optimal the space $\mathcal{B}$ restricts to the set $\mathsf{Vert}(\mathcal{B})$, so any learned knowledge $\bm{\mathrm{\beta}}$ must belong to $\mathsf{Vert}(\mathcal{B})$. This shows that optimal pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ belong to $\mathcal{A}\times\mathsf{Vert}(\mathcal{B})$.
Step (ii). Next, we make use of Assumption˜3.8 and consider only inference functions $\bm{\mathrm{\omega}}$ that satisfy it. We start by considering a pair $(\bm{\mathrm{a}}^{\prime},\bm{\mathrm{b}}^{\prime})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ that attain maximum likelihood, that is, according to Theorem˜3.6:
@@
Now, given $\bm{\mathrm{b}}^{\prime}\in\mathsf{Vert}(\mathcal{B})$, we have to prove that no non-deterministic $\bm{\mathrm{\alpha}}\in\mathcal{A}\setminus\mathsf{Vert}(\mathcal{A})$ can attain maximum likelihood for any of the $\bm{\mathrm{\omega}}^{\prime}\in\Omega$ that correspond to $\bm{\mathrm{b}}^{\prime}$, that is, such that $\bm{\mathrm{\omega}}^{\prime}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})=\bm{\mathrm{b}}^{\prime}(\bm{\mathrm{c}})$ for all $\bm{\mathrm{c}}\in\mathcal{C}$. Since $\mathcal{A}$ is a simplex, we can always construct a non-deterministic conditional probability distribution $\bm{\mathrm{\alpha}}\in\mathcal{A}$ from a convex combination of the vertices $\bm{\mathrm{a}}_{i}\in\mathsf{Vert}(\mathcal{A})$, i.e., for all $\bm{\mathrm{g}}\in\mathcal{G}$ it holds:
@@
where $\lambda\in(0,1)$. When the deterministic JRSs count (Eq.˜6) equals to zero, we know that by Theorem˜3.6 only $(\bm{\mathrm{a}}^{\prime},\bm{\mathrm{b}}^{\prime})$ attains maximum likelihood, whereas $(\bm{\mathrm{a}}^{\prime\prime},\bm{\mathrm{b}}^{\prime})$ is not optimal. Therefore, there exists at least one $\hat{}\bm{\mathrm{g}}\in\mathsf{supp}(p^{*}(\bm{\mathrm{G}}))$ such that
@@
We now have to look at the form of $\bm{\mathrm{f}}$ that can induce such a non-deterministic $\bm{\mathrm{\alpha}}=\lambda\bm{\mathrm{a}}^{\prime}+(1-\lambda)\bm{\mathrm{a}}^{\prime\prime}$. Recalling, the definition of $\bm{\mathrm{\alpha}}$ (Eq.˜2) we have that:
@@
(2) The remaining option is that $\bm{\mathrm{f}}:\mathbb{R}^{n}\to\Delta_{\mathcal{C}}\setminus\mathsf{Vert}(\Delta_{\mathcal{C}})$. In light of this, we rewrite $\bm{\mathrm{f}}$ as follows:
@@
where $p(\bm{\mathrm{c}}\mid\bm{\mathrm{x}}):=\bm{\mathrm{f}}(\bm{\mathrm{x}})_{\bm{\mathrm{c}}}$. Let $\hat{}\bm{\mathrm{c}}^{\prime}:=\bm{\mathrm{a}}^{\prime}(\hat{}\bm{\mathrm{g}})$ and $\hat{}\bm{\mathrm{c}}^{\prime\prime}:=\bm{\mathrm{a}}^{\prime\prime}(\hat{}\bm{\mathrm{g}})$. But $\bm{\mathrm{\alpha}}$ is a convex combination of $\bm{\mathrm{a}}^{\prime}$ and $\bm{\mathrm{a}}^{\prime\prime}$, hence the function $\bm{\mathrm{f}}$ must attribute non-zero probability mass only to the concept vectors $\hat{}\bm{\mathrm{c}}^{\prime}$ and $\hat{}\bm{\mathrm{c}}^{\prime\prime}$. Hence, we rewrite it as:
@@
where $\mathcal{X}$ has measure one. Notice that in this case it holds $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{b}}^{\prime}(\hat{}\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\neq\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{b}}^{\prime}(\hat{}\bm{\mathrm{c}}^{\prime\prime})_{\bm{\mathrm{y}}}$. By Assumption˜3.8 we have that, for any choice of $\lambda\in(0,1)$ it holds that for all $\bm{\mathrm{x}}\in\mathcal{X}$:
@@
where the last step in Eq.˜63 follows from the fact that $\bm{\mathrm{b}}^{\prime}\in\mathsf{Vert}(\mathcal{B})$, so giving point-mass conditional probability distributions.
(1) and (2) together show that any $\bm{\mathrm{\alpha}}\in\mathcal{A}\setminus\mathsf{Vert}(\mathcal{A})$ constructed as a convex combination of $\bm{\mathrm{a}}^{\prime}$ and an $\bm{\mathrm{a}}^{\prime\prime}\neq\bm{\mathrm{a}}^{\prime}$ cannot attain maximum likelihood. Hence, the optimal pairs restrict to $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ and since the count of JRSs (Eq.˜6) is zero, all $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ that are optimal also possess the intended semantics. This concludes the proof. ∎
@@ Appendix D Impact of mitigation strategies on the deterministic JRSs count
We focus on how the count can be updated for some mitigation strategies that we accounted for in Section˜4. We consider those strategies that gives a constraint for the Eq.˜6, namely: multi-task learning, concept supervision, knowledge distillation, and reconstruction.
@@ D.1 Multi-task learning
In the following, we use $\tau$ to indicate an additional learning task with corresponding. Suppose that for a subset $\mathcal{G}^{\tau}\subseteq\mathrm{supp}(p^{*}(\bm{\mathrm{G}}))$, input examples are complemented with additional labels $\bm{\mathrm{y}}^{\tau}$, obtained by applying an inference layer $\bm{\mathrm{\beta}}_{\mathsf{K}^{\tau}}(\bm{\mathrm{g}})$. Here, $\mathsf{K}^{\tau}$ is the conjunction of the prior knowledges for the original and additional task, and it yields a set $\mathcal{Y}^{\tau}$ of additional labels when applied to the examples in $\mathcal{G}^{\tau}$. To understand how multi-task learning impacts the count of JRSs, we have to consider those $\bm{\mathrm{\beta}}:\mathcal{C}\to\mathcal{Y}\times\mathcal{Y}^{\tau}$, where a component $\bm{\mathrm{\beta}}_{\mathcal{Y}}$ maps to the original task labels $\bm{\mathrm{Y}}$ and the other component $\bm{\mathrm{\beta}}_{\mathcal{Y}^{\tau}}$ maps to the augmented labels $\bm{\mathrm{Y}}^{\tau}$. Let $\mathcal{B}^{*}:={\mathcal{B}}\times{\mathcal{B}^{\tau}}$ the space where such functions are defined. As obtained in [19], the constraint for a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B}^{*})$ can be written as:
@@
This term can be readily combined with the one appearing in Theorem˜3.6. Doing so, reduces the number of possible maps $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ that successfully allow a function $\bm{\mathrm{\beta}}\in\mathcal{B}^{*}$ to predict both $(\bm{\mathrm{y}},\bm{\mathrm{y}}^{\tau})$ consistently. Notice that, in the limit where the augmented knowledge comprehends enough additional labels and the support covers the whole , the only admitted solutions consist of:
@@ D.2 Concept supervision
The intuition is the same as in [19]. Assume that we supply concept supervision for a subset of concepts $\bm{\mathrm{G}}_{\mathcal{I}}\subseteq\bm{\mathrm{G}}$, with $\mathcal{I}\subseteq[k]$ and pair the regular log-likelihood objective over the labels with a log-likelihood over the concepts, i.e., $\sum_{i\in\mathcal{I}}\log p_{\theta}(C_{i}=g_{i}\mid\bm{\mathrm{x}})$. Let $\mathcal{G}^{C}\subseteq\mathsf{supp}(p^{*}(\bm{\mathrm{G}}))$ be the subset of values that receive supervision. The only deterministic $\bm{\mathrm{\alpha}}$’s that attain maximum likelihood on the concepts are those that match said supervision, i.e., that satisfy the constraint:
@@
This can be immediately used to obtain an updated (and smaller) JRS count. Note that if $\mathcal{I}=[k]$ and $\mathcal{G}^{C}=\mathcal{G}$, then the only suitable map is $\bm{\mathrm{\alpha}}\equiv\mathrm{id}$. Naturally, this requires dense annotations for all possible inputs, which may be prohibitive.